7) Find the equivalent resistance and the rrent between points A and B if VAB-92Vol 52. 163 232 302
Answer #1

Solution) Resistors are either connected in series or in parallel

Equivalent resistance for series connected resistors is

Rs = R1 + R2 +•••+Rn ( for n resistors in series)

Equivalent resistance for parallel connected resistors is

(1/Rp) = (1/R1) + (1/R2) +•••••+(1/Rn) (for n resistors in parallel)

Res 23 A it、YChis lances 16 lou and su Now Rs Kg y in parallel on th 10. ther epui valent (KS) (3°) Rs 2 Re= (31) (3) 13:24 2.2人 eislau Rey- 38-24